In a titration experiment, an average of 15.2 mL of a 0.25 mol/L NaOH solution was needed to neutralize 20.0 mL of an unknown concentration of HCl. What is the molar concentration of the HCl solution?


1
Expert's answer
2022-06-08T12:09:03-0400

1.

We can find the amount of NaOH by the following formula:

n = CM * V

n(NaOH) = 0,25mol/l * 0,0152 l = 0,0038mol

2.

We find the amount of HCl.

NaOH + HCl = NaCl + H2O

1mol NaOH — 1mol HCl

0,0038mol NaOH — X

X = 0,0038 mol HCl

3.

We find the concentration of HCl by the following formula:

"C_M= \cfrac {n} {V}= \cfrac {0,0038}{0,02}= 0,19mol/l"


Answer: 0,19 mol/l